The length, width, and thickness of the glass are 11cm, 6.45cm, and 0.025cm respectively. If the density of the glass is 2150kgm%B3, find the mass of the glass


1
Expert's answer
2021-04-12T06:55:26-0400

"m=\rho V=\rho \cdot a\cdot b\cdot c="


"=2150\cdot0.11\cdot0.0645\cdot0.00025=0.0038\ (kg)=3.8\ (g)" . Answer